OpenStudy (anonymous):

solve the inequality. graph the solution set and write it in interval notation. -9 \[\le\] -2x + 5 < 1 Is it (7, infty) or the other way around.

OpenStudy (anonymous):

\[-9\le-2x+5<1\]?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

i just started using this so Im still trying to fig out how to use some of the signs...

OpenStudy (anonymous):

\[-9-5\le-2x<1-5\]\[-14\le-2x<-4\]\[\frac{-14}{-2}\le x<\frac{-4}{-2}\]\[14\ge x>2\]\[(2,\infty)\]?

OpenStudy (anonymous):

the last one I was wrong

OpenStudy (anonymous):

wouldnt you divide the 2 over the 14 to get 7 or am I way off...

OpenStudy (anonymous):

I caculator wrong

OpenStudy (anonymous):

\[7\ge x>2\]

OpenStudy (anonymous):

you r rite... I was looking at the graph wrong

OpenStudy (anonymous):

−9−5≤−2x+5-5<1−5 −14≤−2x<−4 −14/2 ≤ -x < −4/2 −7≤-x<−2 \[7 \ge x > 2\] or 2<x≤7 \[ x \in ]2,7] \]

OpenStudy (anonymous):

|x|x>2 , x<=2|

OpenStudy (anonymous):

no infty

OpenStudy (anonymous):

got it.... thank you both very much

OpenStudy (anonymous):

ty